OpenStudy (anonymous):

the factorization x squared -3x +2 =

OpenStudy (anonymous):

!!

OpenStudy (anonymous):

HELP!

OpenStudy (pfenn1):

Sorry, I was confusing myself. I know what the answer is but I don't know how to explain it to you. \[x ^{2} - 3x+2=\]

OpenStudy (pfenn1):

You need to find factors of 2 that add up to -3. Since 2 can be written as the product of 2 and 1 or -1 and -2, and since -1 - 2 = -3, then use -1 and -2. You know from multiplying polynomials that this quadratic is formed from multiplying two factors of the form "(x + m)(x + n)", for some numbers m and n. So \[x ^{2}-3x+2=(x-1)(x_2)\]

OpenStudy (pfenn1):

That should be \[(x-1)(x-2)\]